what is the breaking strength of a rope. is there any formula to calculate it ?

Grade:Upto college level

2 Answers

Rinkoo Gupta
9 years ago
CALCULATION OF APPROXIMATE BREAKING STRENGTH (B.S) AND SAFE WORKING LOAD (S.W.L) FORMANILAROPE
Method of finding theBreaking Strength (B.S)is to divide the square of the diameter of the rope in millimetres by200.

Example of adiameter 24mmManila Rope:
Breaking strength = diameter² / 200
= 24² / 200
= 576 / 200
=2.88 tonnes (approx. 3 tonnes)CALCULATION OF APPROXIMATE BREAKING STRENGTH (B.S) AND SAFE WORKING LOAD (S.W.L) FOR POLYPROPYLENE ROPE

Method of finding theBreaking Strength (B.S)is to divide the square of the diameter of the rope in millimetres by77 tonnes.

Example of adiameter 24mmPolypropylene Rope:

Breaking strength = diameter² / 77
= 24² / 77
= 576 / 77
=7.48 tonnes (approx. 7 tonnes)
ALCULATION OF APPROXIMATE BREAKING STRENGTH (B.S) AND SAFE WORKING LOAD (S.W.L) FOR POLYETHYLENE ROPE

Method of finding theBreaking Strength (B.S)is to divide the square of the diameter of the rope in millimetres by106 tonnes.

Example of adiameter 24mmPolyethylene Rope:

Breaking strength = diameter² / 106
= 24² / 106
= 576 / 106
=5.43 tonnes (approx. 5 tonnes)
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty

Rakesh
5 years ago
what is the breaking strength of a rope. is there any formula to calculate it ?
What is  Breaki strength = diameter² / 77
What is 77